# Sigma: Differential Rescaling of Query, Key and Value for Efficient Language Models

### v1

@@ Abstract
We introduce Sigma, an efficient large language model specialized for the system domain, empowered by a novel architecture including DiffQKV attention, and pre-trained on our meticulously collected system domain data. DiffQKV attention significantly enhances the inference efficiency of Sigma by optimizing the Query (Q), Key (K), and Value (V) components in the attention mechanism differentially, based on their varying impacts on the model performance and efficiency indicators. Specifically, we (1) conduct extensive experiments that demonstrate the model’s varying sensitivity to the compression of K and V components, leading to the development of differentially compressed KV, and (2) propose augmented Q to expand the Q head dimension, which enhances the model’s representation capacity with minimal impacts on the inference speed. Rigorous theoretical and empirical analyses reveal that DiffQKV attention significantly enhances efficiency, achieving up to a 33.36% improvement in inference speed over the conventional grouped-query attention (GQA) in long-context scenarios. We pre-train Sigma on 6T tokens from various sources, including 19.5B system domain data that we carefully collect and 1T tokens of synthesized and rewritten data. In general domains, Sigma achieves comparable performance to other state-of-arts models. In the system domain, we introduce the first comprehensive benchmark AIMicius, where Sigma demonstrates remarkable performance across all tasks, significantly outperforming GPT-4 with an absolute improvement up to 52.5%.
@@ 1 Introduction
In recent years, significant progress has been made in the development of large language models (LLMs), which have demonstrated remarkable performance across a wide range of domains (Bubeck et al., 2023; Jiang et al., 2023a; GLM et al., 2024; Dubey et al., 2024; Yang et al., 2024). Meanwhile, a novel research direction, known as the “system domain”, has emerged with promising potential to further accelerate AI development through automated optimization of AI infrastructure (Xiong et al., 2024; Shi et al., 2024; Hu et al., 2024b). This domain focuses on leveraging AI models to autonomously validate, evaluate, diagnose, and optimize the key components of AI infrastructure (e.g., hardware, configurations, cloud services, databases, and workloads). By advancing the robustness, efficiency, and scalability of these foundational systems, the system domain offers a powerful pathway to enhance the capabilities and performance of AI. Despite its promise, however, this important area of research has yet to receive commensurate attention.
@@
The DiffQKV attention adopted by Sigma substantially improves its inference efficiency by mitigating the bottleneck associated with KV cache (Pope et al., 2023). [^1] KV cache is a common technique adopted by the decoder-only Transformer architecture (Vaswani et al., 2017), which stores the Key (K) and Value (V) vectors in the attention operation for future reuse at each decoding step. It can consume a substantial amount of memory (Pope et al., 2023) and place substantial demands on memory bandwidth (Shazeer, 2019; Ribar et al., 2024). Though there have already been significant efforts to address the KV cache issue (Ainslie et al., 2023; Kwon et al., 2023; Luohe et al., 2024; Zhang et al., 2024b), prior studies tend to treat the compression of K and V vectors uniformly and rarely take the optimization of Q into consideration. In contrast, DiffQKV attention differentially optimizes the Query (Q), Key (K), and Value (V) components in the attention mechanism with tailored strategies, based on their varying impacts on the model performance and efficiency indicators. Specifically, it involves two critical techniques: differentially compressed KV and augmented Q. Differentially compressed KV is grounded on our experimental findings that the overall model performance is more sensitive to compression in V vectors than K, both in terms of the dimension and the number of heads. We leverage this disparity by implementing a more aggressive compression algorithm for K, whereas V employs a lighter form of compression. Despite the relatively low compression on V, we can further optimize this part by loading V vectors selectively during inference. Due to the high sparsity exhibited by the attention scores (Xiao et al., 2023; Zhang et al., 2024b), it can still preserve the model performance with only a small number of V vectors to approximate the results, while largely reducing the memory usage. Augmented Q involves adopting a higher dimension for the Q head compared to the KV heads, as we discover that introducing extra parameters to the Q head components can effectively boost the model performance. With minimal impacts on the inference speed, augmented Q can to some degree counteract the performance decline that inevitably results from KV compression. Rigorous theoretical and empirical analyses reveal that Sigma’s DiffQKV significantly enhances efficiency, achieving up to a 33.36% improvement in inference speed over the conventional grouped-query attention (GQA) in long-context scenarios.
@@ 2 DiffQKV Attention
In the standard Multi-Head Attention (MHA) introduced by Vaswani et al. (2017), queries, keys, and values consistently employ an equal number of heads, with each head maintaining the same dimension. Additionally, existing autoregressive decoding approaches typically require loading the entire KV cache before computing attention scores at each generation step.
Our research explores a more generalized form of attention mechanism, termed attention with differential rescaling of QKV (DiffQKV attention), which involves two crucial variations: (1) the Q, K, and V components can have varying numbers of heads and differing dimensions per head; (2) during inference, the K and V caches are retrieved respectively with distinct strategies. Its inference process is formally defined as follows.
2.1 Formulation of DiffQKV Attention
The arguments for DiffQKV attention include:
  • $h_{t}\in\mathbb{R}^{d_{m}}$: the attention input for the $t$-th token within a specific attention layer.
  • $n_{q}^{h},n_{k}^{h},n_{v}^{h}$: number of the Q, K, and V heads, respectively. Typically, the number of Q heads is set as a multiple of K heads and V heads. This can be expressed as $n_{q}^{h}=g_{k}\cdot n_{k}^{h}=g_{v}\cdot n_{v}^{h}$, where $g_{k}$ and $g_{v}$ are positive integers.
  • $d_{q}^{h},d_{k}^{h},d_{v}^{h}$: dimension of the Q, K, and V heads, respectively.
  • $W_{Q},W_{K},W_{V},W_{O}$: weight tensors for query, key, value, and output projections, with shapes of $[n_{q}^{h},d_{m},d_{q}^{h}],[n_{k}^{h},d_{m},d_{k}^{h}],[n_{v}^{h},d_{m},d_{v}^{h}]$, and $[n_{v}^{h}*d^{h}_{v},d_{m}]$, respectively.
  • &$\&K_{\text{cache}},\&V_{\text{cache}}$: memory addresses for the K and V caches, respectively.
We begin by transforming the input $h_{t}$ to Q, K, V heads through:
$q_{t}$ $=h_{t}\cdot W_{Q},$ (1)
$k_{t}$ $=h_{t}\cdot W_{K},$ (2)
$v_{t}$ $=h_{t}\cdot W_{V},$ (3)
where the dimensions of $q_{t}$, $k_{t}$, and $v_{t}$ are $[n_{q}^{h},d_{q}^{h}]$, $[n_{k}^{h},d_{k}^{h}]$, and $[n_{v}^{h},d_{v}^{h}]$ respectively. Then, we load the previous K vectors from the cache and concatenate it with $k_{t}$.
$K_{\text{t-1}}$ $=\text{\emph{LoadCacheK}}(\&K_{\text{cache}}),$ (4)
$K_{t}$ $=[K_{\text{t-1}};k_{t}],$ (5)
@@ 2.1 Formulation of DiffQKV Attention
where the dimension of $K_{t}$ is $[t,n_{k}^{h},d_{k}^{h}]$. To prepare for the attention score computation, $K_{t}$ and $q_{t}$ are sliced into a series of heads. In addition, since the number of K heads and Q heads can be inconsistent, we employ a GroupSharing operation on $K_{t}$, which allows multiple Q heads to be jointly associated with the same K head:
$[K_{(t,1)};K_{(t,2)};..;K_{(t,n^{h}_{q})}]$ $=\text{\emph{GroupSharing}}(K_{t},n^{h}_{q},n^{h}_{k}),$ (6)
$[{q}_{(t,1)};{q}_{(t,2)};..;{q}_{(t,n^{h}_{q})}]$ $=q_{t},$ (7)
@@
where $q_{(t,i)}\in\mathbb{R}^{d_{k}^{h}}$ represents the $i$-th Q head $(i=1,2,..,n^{h}_{q})$. $K_{(t,i)}$ is in the shape of $[t,d_{k}^{h}]$. The attention scores are then computed as:
$\alpha_{(t,i)}$ $=\text{Softmax}_{i}[\frac{\text{\emph{Attend}}({q}^{\;T}_{(t,i)},K_{(t,i)})}{\sqrt{d_{k}}}],$ (8)
where $\alpha_{(t,i)}\in\mathbb{R}^{t}$. Alternatively, FlexHeadFA can be used to perform attention computations without the inefficient group sharing operation. For more details, we kindly refer readers to § 3.2.
@@
In standard MHA, the Attend function computes the inner products of the Q heads and K heads. Nonetheless, in our paradigm, which accommodates varying dimensions for the Q and K heads, an alternative implementation of the Attend function is necessitated. In our experiments, we implement this part by transforming the dimension of K to the same as V through another feed-forward layer.
Next, we load V cache and prepare the V vectors for computing the outputs.
$V_{\text{t-1}}$ $=\text{\emph{LoadCacheV}}(\&V_{\text{cache}},\alpha_{(t,i)}),$ (9)
$V_{t}$ $=[V_{\text{t-1}};v_{t}],$ (10)
$[V_{(t,1)};V_{(t,2)};..;V_{(t,n^{h}_{q})}]$ $=\text{\emph{GroupSharing}}(V_{t},n^{h}_{q},n^{h}_{v}),$ (11)

### v2

@@ Abstract
We introduce Sigma, an efficient large language model specialized for the system domain, empowered DiffQKV attention, and pre-trained on self-collected system domain data. Given the varing impacts on the model performance and efficienct indicators of Query (Q), Key (K), and Value (V), Sigma use DiffQKV attention to optimize them differentially and significantly enhance inference efficiency. Specifically, we (1) differentially compress K and V, leveraging the model’s varying sensitivity to KV compression, as demonstrated in our extensive experiments, and (2) propose augmented Q to expand the dimension of the Q head, which improves the representation capacity of the model with minimal impacts on the inference speed. Rigorous theoretical and empirical analyses reveal that DiffQKV attention significantly enhances inference efficiency, achieving up to a 33.36% improvement over the conventional grouped-query attention (GQA) in long-context scenarios. We pre-train Sigma on 6T tokens from various sources, including 19.5B system domain data that we carefully collect and 1T tokens of synthesized and rewritten data. In general domains, Sigma performs on par with state-of-the-art models. In the system domain, it excels on AIMicius, the first comprehensive system-domain benchmark we propose, surpassing GPT-4 by up to 52.5% across all tasks.
Machine Learning, ICML
@@ 1 Introduction
In recent years, significant progress has been made in the development of large language models (LLMs), which have demonstrated remarkable performance across a wide range of domains (Bubeck et al., 2023; Jiang et al., 2023a; GLM et al., 2024; Dubey et al., 2024; Yang et al., 2024). Meanwhile, a novel research direction, known as the “system domain,” has emerged with promising potential to further accelerate AI development through automated optimization of AI infrastructure (Xiong et al., 2024; Shi et al., 2024; Hu et al., 2024b). This domain focuses on leveraging AI models to autonomously validate, evaluate, diagnose, and optimize the key components of AI infrastructure (e.g., hardware, configurations, cloud services, databases, and workloads). Despite its promise, however, this important area of research has yet to receive commensurate attention.
@@
The DiffQKV attention substantially improves Sigma’s inference efficiency by mitigating the bottleneck associated with KV cache (Pope et al., 2023). [^1] KV cache is a common technique in the decoder-only Transformer architecture (Radford et al., 2019), which stores Key and Value vectors for future reuse in decoding. It can consume considerable GPU memory (Pope et al., 2023) and place substantial demands on memory bandwidth (Shazeer, 2019; Ribar et al., 2024). Despite significant efforts to address the KV cache issue (Ainslie et al., 2023; Kwon et al., 2023; Luohe et al., 2024; Zhang et al., 2024b), previous studies tend to treat the compression of K and V vectors uniformly and rarely take into account the optimization of Q. In contrast, DiffQKV attention differentially optimizes the Query (Q), Key (K), and Value (V) components in the attention mechanism with tailored strategies, based on their varying impacts on the model performance and efficiency indicators. Specifically, it involves two critical techniques: differentially compressed KV and augmented Q. Differentially compressed KV aggressively compresses K while lightly compressing V, grounded on our experimental findings that model performance is significantly more sensitive to compression in V vectors than in K, both in terms of dimension and the number of heads. Augmented Q adopts a higher dimension for the Q head compared to KV heads, as we discover that introducing extra parameters to Q can effectively boost the model performance. With minimal impacts on the inference speed, augmented Q can to some degree counteract the performance decline that inevitably results from KV compression. Rigorous theoretical and empirical analyses reveal that Sigma’s DiffQKV significantly enhances efficiency, achieving up to a 33.36% improvement in inference speed over the conventional grouped-query attention (GQA) in long-context scenarios.
@@ 2 DiffQKV Attention
In the standard Multi-Head Attention (MHA) introduced by Vaswani et al. (2017), Q, K, and V components consistently employ an equal number of heads, with each head maintaining the same dimension. Our research explores a more generalized form of attention mechanism, termed DiffQKV attention, where QKV can possess distinct numbers of heads as well as different dimensions per head. In this section, we will first formally describe the DiffQKV attention process and then discuss the optimal configurations for it to effectively balance efficiency and model performance.
Formulation of DiffQKV Attention. Let the numbers of QKV heads as $n_{q}^{h},n_{k}^{h},n_{v}^{h}$, respectively, and the dimensions of QKV heads as $d_{q}^{h},d_{k}^{h},d_{v}^{h}$. Formally, given the attention input $h_{t}$, we begin by transforming it to:
$q_{t}=h_{t}\cdot W_{Q},\quad k_{t}=h_{t}\cdot W_{K},\quad v_{t}=h_{t}\cdot W_{V},$
@@
where the dimensions of $q_{t}$, $k_{t}$, $v_{t}$ are $[n_{q}^{h},d_{q}^{h}]$, $[n_{k}^{h},d_{k}^{h}]$, $[n_{v}^{h},d_{v}^{h}]$ respectively. Then, we load the previous K vectors from the cache and concatenate it with $k_{t}$, thereby obtaining $K_{t}$, which has a dimension of $[t,n_{k}^{h},d_{k}^{h}]$. To prepare for the attention score computation, $K_{t}$ and $q_{t}$ are sliced into a series of heads. Since the number of K heads and Q heads can be inconsistent, we employ a GroupSharing operation on $K_{t}$, which allows multiple Q heads to be jointly associated with the same K head (see Appendix, Figure 3):
$[K_{(t,1)};K_{(t,2)};..;K_{(t,n^{h}_{q})}]$ $=\text{\emph{GroupSharing}}(K_{t},n^{h}_{q},n^{h}_{k}),$
$[{q}_{(t,1)};{q}_{(t,2)};..;{q}_{(t,n^{h}_{q})}]$ $=q_{t},$
@@
where $q_{(t,i)}\in\mathbb{R}^{d_{k}^{h}}$ represents the $i$-th Q head. $K_{(t,i)}$ is in the shape of $[t,d_{k}^{h}]$. The attention scores are computed as:
$\alpha_{(t,i)}=\text{Softmax}_{i}[\frac{\text{\emph{Attend}}({q}^{\;T}_{(t,i)},K_{(t,i)})}{\sqrt{d_{k}}}],$
@@
where $\alpha_{(t,i)}\in\mathbb{R}^{t}$. In standard MHA, the Attend function computes the inner products of the Q heads and K heads. Nonetheless, as DiffQKV accommodates varying dimensions for the Q and K heads, an alternative implementation of the Attend function is necessitated. In our experiments, we implement this part by transforming the dimension of K to the same as V through another feed-forward layer. Next, we load previous V cache and compute the outputs as:
$V_{t}=[V_{\text{t-1}};v_{t}],\quad V_{\text{t-1}}$ $=\text{\emph{LoadCacheV}}(\&V_{\text{cache}},\alpha_{(t,i)}),$
$[V_{(t,1)};V_{(t,2)};..;V_{(t,n^{h}_{q})}]$ $=\text{\emph{GroupSharing}}(V_{t},n^{h}_{q},n^{h}_{v}),$
$o_{(t,i)}$ $=\sum\nolimits_{j=1}^{t}\alpha_{(j,i)}V_{(j,i)},$
$x_{t}$ $=W_{O}[o_{(t,1)};o_{(t,2)};...;o_{(t,n_{v}^{h})}].$